# Fact or Fiction: Can Truthful Mechanisms Eliminate Federated Free Riding?

### v2

@@ 4 Eliminating Free Riding via Penalization
$\lambda_{i}:=\frac{m^{*}(\sum\bm{m})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m}_{-i}+m^{*})^{2}}\bigg{)}^{2}.$ (6)
@@
$\Delta\ell_{i}:=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m}_{-i}+m^{*})}\bigg{)}.$ (7)
@@
3:  Server receives contract payment from each agent: $\Delta\ell_{i}\leftarrow\frac{\alpha}{4}\big{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m_{i}(\sum\bm{m})}\big{)}$
4:  Run Penalized Federated Learning: $\bm{w}^{T}\leftarrow$ PFL($\bm{m},\bm{c},h,T,E,\bm{w}^{1},f,\gamma,L,\sigma^{2},\alpha$)
5:  Compute each agent reward $r_{i}$ via the cost truthfulness game in Equation (10)
@@ 5 Fact: Eliminating Free-Riding With Untruthful Agents
$P_{ct}(c):=\begin{cases}\Delta\ell_{i}-\frac{3}{n}\sum\limits_{j\neq i\in[n]}\Delta\ell_{j}\;&\text{if }C_{a}<c<C_{b},\\ \Delta\ell_{i}\;&\text{else}.\end{cases}\;C_{a},C_{b}\sim f_{C}\text{ randomlysampled costs}$ (10)
@@
The probability of an agent “winning” in the truthfulness mechanism in Equation (10), given a reported cost $c$, is $\upsilon(c):=2F_{C}(c)(1-F_{C}(c))$, where $F_{C}$ is the CDF of $f_{C}$. The crux of the truthfulness competition is that each agent $i$ never receives its own reward $\Delta\ell_{i}$. Instead, as shown in Equation (10), each agent pays $\Delta\ell_{i}$ to the server before training regardless of the mechanism outcome. After training, if the agent “wins” the competition, by having a cost sandwiched between $C_{a}$ and $C_{b}$, then it receives triple the average of all other agent rewards $\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}$. The average reward is tripled, since two other agents will lose the competition and receive no reward. By making the reward decoupled from an agent’s own reward $\Delta\ell_{i}$, each agent can no longer increase or decrease its reward by altering how much data $m_{i}$ they use for federated training. Now, they can only affect the likelihood of winning the competition by choosing what cost $c_{i}$ to report to the server.
@@
$\ell_{i,Fact}(m_{i},c)=\underbrace{\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}+c_{i}m_{i}+P_{fr}(m_{i},c)}_{\ell_{i,PFL}(m_{i},c)}+P_{ct}(c).$ (11)
@@
[Individually Rational] When using their best strategy, agents participating in Fact either (i) receive loss equivalent to local training if they lose the competition or (ii) reduced loss if they win (shown in Equation 10). Thus, Fact is IR as agents using their best strategy never receive worse loss than if they did not participate. In fact, agents receive lower loss in expectation.
@@ Appendix C Proofs
$\lambda_{i}:=\frac{m^{*}(\sum\bm{m})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m}_{-i}+m^{*})^{2}}\bigg{)}^{2}.$ (17)
@@
$\Delta\ell_{i}:=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m}_{-i}+m^{*})}\bigg{)}.$ (18)
@@
$\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})$ $=\frac{\gamma\sigma^{2}L}{2}\big{(}\frac{1}{m^{*}}-\frac{1}{m^{*}+\sum\bm{m}_{-i}}\big{)}-\lambda_{i}(\frac{c_{i}}{2\lambda_{i}}-\frac{\gamma\sigma^{2}L}{4\lambda_{i}(\sqrt{\frac{\gamma\sigma^{2}L}{2c_{i}}}+\sum\bm{m}_{-i})^{2}})^{2},$
@@
$\frac{\gamma\sigma^{2}L}{2}\frac{\sum\bm{m}_{-i}}{m^{*}(m^{*}+\sum\bm{m}_{-i})}-\frac{1}{4\lambda_{i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})^{2}}\bigg{)}^{2}\leq 0,$
@@
$\lambda_{i}:=\frac{m_{i}(\sum\bm{m})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m})^{2}}\bigg{)}^{2}.$ (21)
@@
$\Delta\ell_{i}:=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m_{-i}}+m^{*})}\bigg{)}.$ (22)
@@
$=\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}+c_{i}m_{i}+P_{fr}(m_{i},c)+\Delta\ell_{i}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}$ (26)
Taking the partial derivative with respect to $m_{i}$ and setting it equal to zero results in,
$\frac{\partial}{\partial m_{i}}\mathbb{E}_{f_{C}}[\ell_{i,Fact}(m_{i},c)]=\frac{\partial}{\partial m_{i}}\bigg{[}\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}+c_{i}m_{i}+P_{fr}(m_{i},c)\bigg{]}=0$ (27)
From Theorem 5, we know that this results in $m_{i}^{*}=\sqrt{\frac{\gamma\sigma^{2}L}{2c}}$. Now, when taking the partial derivative with respect to $c$ we can plug in $m_{i}^{*}$ (to ensure the critical point is evaluated as zero),
$\frac{\partial}{\partial c}\bigg{[}\underbrace{\frac{\gamma\sigma^{2}L}{2(m_{i}^{*}+\sum\bm{m}_{-i})}+c_{i}m_{i}^{*}+P_{fr}(m_{i}^{*},c)+\Delta\ell_{i}}_{\ell_{i,l}(m_{i}^{*})}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}=0.$ (28)
By Lemma 1 and specifically Equation (7), we find that the first four terms above are equivalent to the local agent loss at its local optimum. Therefore, we can rewrite the equation above as,
$\frac{\partial}{\partial c}\bigg{[}\underbrace{\frac{\gamma\sigma^{2}L}{2m_{i}^{*}}+c_{i}m_{i}^{*}}_{\ell_{i,l}(m_{i}^{*})}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}$ $=0\longrightarrow-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}\upsilon(c)\bigg{]}=0,$
$-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}2F_{C}(c)(1-F_{C}(c))\bigg{]}$ $=0.$ (29)
@@
The final line follows from Lemma 5. Since $F_{C}(c)$ is a CDF, its range is $[0,1]$. The function $\upsilon(c)=2F_{C}(c)(1-F_{C}(c))$ achieves is global maximum when $F_{C}(c)=1/2$ (its critical point). Therefore, the value of $c$ which satisfies Equation (29) is one such that $F_{C}(c)=1/2$. By Assumption 5, we know that the median value is assumed by each agent $i$ as $c_{i}$, since each agent believes its cost is equally likely to be larger or smaller than any other agent’s cost. Thus, $F_{C}(c_{i})=1/2$ and $c^{*}=c_{i}$. Plugging this optimal value of $c$ back into the optimal value for $m_{i}$ leads to the following optimum: $(m_{i},c)^{*}=(\sqrt{\frac{\gamma\sigma^{2}L}{2c_{i}}},c_{i})$.

### v3

@@ 4 Eliminating Free Riding via Penalization
$\lambda_{i}:=\frac{m^{*}(\sum\bm{m}_{-i}+m^{*})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m}_{-i}+m^{*})^{2}}\bigg{)}^{2}.$ (6)
@@
$\Delta\ell_{i}(m^{*}):=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m}_{-i}+m^{*})}\bigg{)}.$ (7)
@@
3:  Server receives contract fees: $\Delta\ell_{i}(m_{i},c_{i})\leftarrow\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{2m_{i}(m_{i}+\sum\bm{m}_{-i})}-P_{fr}(m_{i},c_{i})$.
4:  $\triangleright$ ($P_{fr}$ cancels out in PFL, so $\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{2m_{i}(m_{i}+\sum\bm{m}_{-i})}$ is the effective contract fee).
5:  Run Penalized Federated Learning: $\bm{w}^{T}\leftarrow$ PFL($\bm{m},\bm{c},h,T,E,\bm{w}^{1},f,\gamma,L,\sigma^{2},\alpha$).
6:  Compute each agent reward $r_{i}$ via the cost truthfulness game in Equation (10).
@@ 5 Fact: Eliminating Free-Riding With Untruthful Agents
$P_{ct}(m_{i},c):=\begin{cases}\Delta\ell_{i}(m_{i},c)-\frac{3}{n}\sum\limits_{j\neq i\in[n]}\Delta\ell_{j}\;&\text{if }C_{a}<c<C_{b},\\ \Delta\ell_{i}(m_{i},c)\;&\text{else}.\end{cases}\;C_{a},C_{b}\sim f_{C}\text{randomly sampled costs}$
@@
The probability of an agent “winning” in the truthfulness mechanism in Equation (10), given a reported cost $c$, is $\upsilon(c):=2F_{C}(c)(1-F_{C}(c))$, where $F_{C}$ is the CDF of $f_{C}$. Similar to Equation (7), the multi-variate $\Delta\ell_{i}$ is defined as $\Delta\ell_{i}(m_{i},c):=\ell_{i,l}(m_{i},c)-\ell_{i,PFL}(m_{i},c)$. Since $\Delta\ell_{i}$ is computed by the server, it uses $c$ instead of $c_{i}$ for the local loss $\ell_{i,l}(m_{i},c)$. The crux of the truthfulness competition is that each agent $i$ never receives its own reward $\Delta\ell_{i}$. Instead, as shown in Equation (10), each agent pays $\Delta\ell_{i}$ to the server before training regardless of the mechanism outcome. After training, if the agent “wins” the competition, by having a cost sandwiched between $C_{a}$ and $C_{b}$, then it receives triple the average of all other agent rewards $\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}$. The average reward is tripled, since two other agents will lose the competition and receive no reward. By making the reward decoupled from an agent’s own reward $\Delta\ell_{i}$, each agent can no longer increase or decrease its reward by altering how much data $m_{i}$ they use for federated training. Now, they can only affect the likelihood of winning the competition by choosing what cost $c_{i}$ to report to the server.
@@
$\ell_{i,Fact}(m_{i},c)=\underbrace{\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}+c_{i}m_{i}+P_{fr}(m_{i},c)}_{\ell_{i,PFL}(m_{i},c)}+P_{ct}(m_{i},c).$ (11)
@@
{corollary}
[Individually Rational] When using their best strategy, agents participating in Fact will never receive worse loss than training alone, $\ell_{i,Fact}(m_{i}^{*},c_{i}^{*})\leq\ell_{i,l}(m_{i,l}^{*})$.
@@
[Individually Rational] When using their best strategy, agents participating in Fact either (i) receive loss equivalent to local training if they lose the competition or (ii) reduced loss if they win (shown in Equation 10). In fact, agents receive lower loss in expectation.
@@ Appendix C Proofs
$\lambda_{i}:=\frac{m^{*}(\sum\bm{m}_{-i}+m^{*})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m}_{-i}+m^{*})^{2}}\bigg{)}^{2}.$ (17)
@@
$\Delta\ell_{i}(m^{*}):=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m}_{-i}+m^{*})}\bigg{)}.$ (18)
@@
$\frac{\gamma\sigma^{2}L}{2}\frac{\sum\bm{m}_{-i}}{m^{*}(m^{*}+\sum\bm{m}_{-i})}-\frac{1}{4\lambda_{i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})^{2}}\bigg{)}^{2}\geq 0,$
@@
$\lambda_{i}:=\frac{m_{i}(\sum\bm{m}_{-i}+m^{*})}{(2-\alpha)\gamma\sigma^{2}L\sum\bm{m}_{-i}}\bigg{(}c_{i}-\frac{\gamma\sigma^{2}L}{2(\sum\bm{m}_{-i}+m^{*})^{2}}\bigg{)}^{2}.$ (21)
@@
$\Delta\ell_{i}(m^{*}):=\ell_{i,l}(m^{*})-\ell_{i,PFL}(m^{*})=\frac{\alpha}{4}\bigg{(}\frac{\gamma\sigma^{2}L\sum\bm{m}_{-i}}{m^{*}(\sum\bm{m_{-i}}+m^{*})}\bigg{)}.$ (22)
@@
$=\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}+c_{i}m_{i}+P_{fr}(m_{i},c)+\Delta\ell_{i}(m_{i},c)-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}.$ (26)
The term $\Delta\ell_{i}$ is defined in Lemma 1, specifically Equation 7. It is the contract fee paid to the server, and thus is also a function of the reported cost $c$,
$\Delta\ell_{i}(m_{i},c)$ $=\ell_{i,l}(m_{i},c)-\ell_{i,PFL}(m_{i},c),$
$=\frac{\gamma\sigma^{2}L}{2m_{i}}+cm_{i}-\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}-cm_{i}-P_{fr}(m_{i},c),$
$=\frac{\gamma\sigma^{2}L}{2m_{i}}-\frac{\gamma\sigma^{2}L}{2(m_{i}+\sum\bm{m}_{-i})}-P_{fr}(m_{i},c).$ (27)
Substituting this back into Equation (26) yields,
$\frac{\gamma\sigma^{2}L}{2m_{i}}+c_{i}m_{i}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}.$ (28)
The partial derivative of $\mathbb{E}_{f_{C}}[\ell_{i,Fact}(m_{i},c)]$ with respect to $c$ yields,
$\frac{\partial}{\partial c}\bigg{[}\frac{\gamma\sigma^{2}L}{2m_{i}}+c_{i}m_{i}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}=0,$
$\frac{\partial}{\partial c}\bigg{[}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}=0,$
$-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}\upsilon(c)\bigg{]}=-\frac{3}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\frac{\partial}{\partial c}\bigg{[}2F_{C}(c)(1-F_{C}(c))\bigg{]}=0.$ (29)
@@
The final line follows from Lemma 5. Since $F_{C}(c)$ is a CDF, its range is $[0,1]$. The function $\upsilon(c)=2F_{C}(c)(1-F_{C}(c))$ achieves is global maximum when $F_{C}(c)=1/2$ (its critical point). Therefore, the value of $c$ which satisfies Equation (29) is one such that $F_{C}(c)=1/2$. By Assumption 5, the median value is assumed by each agent $i$ as $c_{i}$, since each agent believes its cost is equally likely to be larger or smaller than any others’. Thus, $F_{C}(c_{i})=1/2$ and $c^{*}=c_{i}$. The partial derivative of $\mathbb{E}_{f_{C}}[\ell_{i,Fact}(m_{i},c)]$ with respect to $m_{i}$ yields,
$\frac{\partial}{\partial m_{i}}\bigg{[}\frac{\gamma\sigma^{2}L}{2m_{i}}+c_{i}m_{i}-\frac{3\upsilon(c)}{n}\sum_{j\neq i\in[n]}\Delta\ell_{j}\bigg{]}=0,$
$c_{i}-\frac{\gamma\sigma^{2}L}{2m_{i}^{2}}=0,$
$\frac{\gamma\sigma^{2}L}{c_{i}}=2m_{i}^{2}\longrightarrow m_{i}=\sqrt{\frac{\gamma\sigma^{2}L}{2c_{i}}}.$ (30)
Thus, the optimal value for $m_{i}$ is $m_{i}^{*}=\sqrt{\frac{\gamma\sigma^{2}L}{2c_{i}}}$. Plugging this optimal value of $c$ back into the optimal value for $m_{i}$ leads to the following optimum: $(m_{i},c)^{*}=(\sqrt{\frac{\gamma\sigma^{2}L}{2c_{i}}},c_{i})$.
Corollary 1 (Assurance of IR at Optimum).
When using their best strategy, agents participating in Fact will never receive worse loss than training alone, $\ell_{i,Fact}(m_{i}^{*},c_{i}^{*})\leq\ell_{i,l}(m_{i,l}^{*})$.
Proof.
First, it is important to note that the optimal data contributions in Fact and locally are equivalent, $m_{i,l}^{*}=m_{i}^{*}$, as shown in Theorems 2 and 1. Thus, both are denoted below as simply $m^{*}$. Similarly, the optimal cost $c_{i}^{*}$ in Fact is simply the true cost $c_{i}$. This too will be used below instead of $c_{i}^{*}$ Now, to prove this Lemma, it must be true that $\ell_{i,Fact}(m^{*},c_{i})-\ell_{i,l}(m^{*})\leq 0$,
$\ell_{i,Fact}(m^{*},c_{i})-\ell_{i,l}(m^{*})$
$=\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+c_{i}m^{*}+P_{fr}(m^{*},c_{i})+P_{ct}(m^{*},c_{i})-(\frac{\gamma\sigma^{2}L}{2m^{*}}+c_{i}m^{*}),$
$=\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+P_{fr}(m^{*},c_{i})+P_{ct}(m^{*},c_{i})-\frac{\gamma\sigma^{2}L}{2m^{*}},$
$\leq\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+P_{fr}(m^{*},c_{i})+\Delta\ell_{i}(m^{*},c_{i})-\frac{\gamma\sigma^{2}L}{2m^{*}},$
$=\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+P_{fr}(m^{*},c_{i})+\ell_{i,l}(m^{*},c_{i})-\ell_{i,PFL}(m^{*},c_{i})-\frac{\gamma\sigma^{2}L}{2m^{*}},$
$=\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+\ell_{i,l}(m^{*},c_{i})-\big{(}\frac{\gamma\sigma^{2}L}{2(m^{*}+\sum\bm{m}_{-i})}+c_{i}m^{*}\big{)}-\frac{\gamma\sigma^{2}L}{2m^{*}},$
$=0.$ (31)